# A Novel State Space Model with Local Enhancement and State Sharing for Image Fusion

### v1

@@ Abstract.
In image fusion tasks, images from different sources possess distinct characteristics. This has driven the development of numerous methods to explore better ways of fusing them while preserving their respective characteristics. Mamba, as a state space model, has emerged in the field of natural language processing. Recently, many studies have attempted to extend Mamba to vision tasks. However, due to the nature of images different from casual language sequences, the limited state capacity of Mamba weakens its ability to model image information. Additionally, the sequence modeling ability of Mamba is only capable of spatial information and cannot effectively capture the rich spectral information in images. Motivated by these challenges, we customize and improve the vision Mamba network designed for the image fusion task. Specifically, we propose the local-enhanced vision Mamba block, dubbed as LEVM. The LEVM block can improve local information perception of the network and simultaneously learn local and global spatial information. Furthermore, we propose the state sharing technique to enhance spatial details and integrate spatial and spectral information. Finally, the overall network is a multi-scale structure based on vision Mamba, called LE-Mamba. Extensive experiments show the proposed methods achieve state-of-the-art results on multispectral pansharpening and multispectral and hyperspectral image fusion datasets, and demonstrate the effectiveness of the proposed approach. Code will be made available.
Multispectral pansharpening, multispectral and hyperspectral image fusion, Mamba, local-enhanced network
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Reconstruction
@@ 1. Introduction
In addressing the first question, two pioneering works stand out: Vim (Zhu et al., 2024), which introduced a bi-directional vision Mamba block tailored for image processing, and VMamba (Liu et al., 2024a), which proposed a scanning mechanism for images to tackle the causality issue inherent in Mamba’s unsuitability for vision tasks. As for the rest of the issues, there has been a lack of exploration of Mamba in the image fusion field. In this paper, we propose two bespoke techniques named local-enhanced vision Mamba (LEVM) block and state sharing to address issues 2) and 3), as its conception shown in Fig. 2, respectively enhancing Mamba’s perception of local information and fully exploiting spatial and spectral information in the SSM’s state.
The proposed method is more suitable for image fusion tasks with two advantages to the proposed methods. The first advantage is that the proposed LEVM block can represent local and global spatial information. Local information brings rich spatial details, while global information injects long-range pixel dependencies, thus enhancing the detail recovery in the fusion process. The second advantage is that the state sharing method allows information to be shared between layers in the adjacent flow and skip-connected flow. Furthermore, we have also proposed the spatial-spectral learning (S2L) of SSM for the state sharing method. The S2L maps the input image to the state space for spatial representation, and considers the state space as the basis for spectral representation, thereby achieving interaction between spatial and spectral information. To the best of our knowledge, there have been no improvements made to vision Mamba from the perspective of spatial and spectral learning specifically for image fusion tasks. Finally, we propose the LE-Mamba, which is constructed based on the U-Net (Ronneberger et al., 2015a) baseline with the inclusion of the LEVM block and the state sharing technique. The proposed LE-Mamba can achieve superior fusion performance.
@@
The rest of the paper will be organized as follows: In Sect. 2, we review the related work on SSMs, then introduce deep learning-based methods on multispectral pansharpening, as well as multispectral and hyperspectral image fusion tasks. In Sect. 3, we revisit the background knowledge of SSM. Then, in Sect. 4, we elaborate the proposed method with the overall architecture, the local-enhanced vision Mamba block, and the state sharing technique in Sect. 4.1, Sect. 4.2 and Sect. 4.3, respectively. Afterward, Sect. 4.4 analyzes the computational complexity of various common neural network operators. At last, we conduct extensive experiments and ablation studies on the proposed architecture to validate its effectiveness in Sect. 5.
@@ 2.2. Deep Learning Methods for Image Fusion
For the multispectral pansharpening task, many CNN-based methods have emerged and obtained promising fused images, including DiCNN (He et al., 2019), PanNet (Yang et al., 2017), and FusionNet (Deng et al., 2020). However, the local feature representation obtained by CNNs hinders better fusion results. For the multispectral and hyperspectral image fusion, there are also many outstanding DL-based works including InvFormer (Zhou et al., 2022b) and MiMO-SST (Fang et al., 2024). However, these methods bear a significant computational burden due to the quadratic complexity of the Transformer. Despite many efforts to mitigate this issue (Han et al., 2023; Wang et al., 2021; Liu et al., 2021), these methods usually lead to a performance drop.
@@ 4.1. Overall Architecture
In some high-level tasks (e.g., classification (Liu et al., 2024a; Zhu et al., 2024) and segmentation (Liu et al., 2024b; Ma et al., 2024; Ruan and Xiang, 2024; Liu et al., 2024a)), architectures similar to Meta-formers (Yu et al., 2022) are commonly used, where plain backbones can offer lower parameter counts, smaller computational loads, and lower latency. However, for the image fusion task, we opted for a multi-scale architecture akin to U-Net (Ronneberger et al., 2015b), which has been proven effective (Cao et al., 2024). The designed network is illustrated in Fig. 2(d), featuring an encoder-decoder architecture. The input comprises LRMS and PAN images, while the output is fused images. The encoder and decoder are composed of multiple LEVM layers, with each LEVM layer comprising several LEVM blocks.
@@ 5.4.3. LEVM Outperforms Previous Attentions
To compare the proposed LEVM with previously common Attention mechanisms, we select common attention and its variants including traditional Attention (Dosovitskiy et al., 2021), PVT Attention (Wang et al., 2021), Swin Attention (Liu et al., 2021), and Linear Attention (Han et al., 2023). We replace the LEVM block with the corresponding Attention block (including FFN) and train them uniformly on the WV3 dataset until convergence. Their fusion performance is shown in Tab. 6. It can be seen that the traditional Attention performs the worst, which is because using Attention in the first layer of the network occupies an extremely large GPU memory, leading to the inability to train. Therefore, most methods based on traditional Attention perform patch embedding in the first layer to downsample the image size, which is disastrous for tasks like image fusion that require low-level information. The performance of Linear Attention is also relatively poor, which is caused by the approximation errors introduced by its approximation of the Attention mechanism. Next are PVT Attention and Swin Attention, both of which impose restrictions on Attention. PVT downsamples both the keys and values, resulting in information loss, while Swin limits the global information to the local regions. The proposed LEVM outperforms previous attention operations.
Table 6. Results of different self-attention (attention) types on WV3 dataset.
@@
5.4.4. Enlarging SSM State Capacity Does not Solve Issue 2)
@@ 6. Conclusion
Based on the observation that sequence modeling in the Mamba network doesn’t fully consider the characteristics of the image fusion task, this paper designs a multi-scale network architecture equipped with the local-enhanced vision Mamba (LEVM) block and the state sharing technique, called LE-Mamba. The LEVM mitigates the information loss caused by limited states in the Mamba network by combining local and global information to enhance fused image details. Additionally, the state sharing technique shares the state between layers in adjacent and skip-connected flows, enabling the deep layers of the network to contain richer spatial information. Then, the fused image exhibits improved spatial and spectral details by incorporating spatial-spectral learning (S2L). The proposed network architecture achieves state-of-the-art performance in widely used image fusion datasets.

### v2

@@ Abstract.
In image fusion tasks, images from different sources possess distinct characteristics. This has driven the development of numerous methods to explore better ways of fusing them while preserving their respective characteristics. Mamba, as a state space model, has emerged in the field of natural language processing. Recently, many studies have attempted to extend Mamba to vision tasks. However, due to the nature of images different from causal language sequences, the limited state capacity of Mamba weakens its ability to model image information. Additionally, the sequence modeling ability of Mamba is only capable of spatial information and cannot effectively capture the rich spectral information in images. Motivated by these challenges, we customize and improve the vision Mamba network designed for the image fusion task. Specifically, we propose the local-enhanced vision Mamba block, dubbed as LEVM. The LEVM block can improve local information perception of the network and simultaneously learn local and global spatial information. Furthermore, we propose the state sharing technique to enhance spatial details and integrate spatial and spectral information. Finally, the overall network is a multi-scale structure based on vision Mamba, called LE-Mamba. Extensive experiments show the proposed methods achieve state-of-the-art results on multispectral pansharpening and multispectral and hyperspectral image fusion datasets, and demonstrate the effectiveness of the proposed approach. Code can be accessed at https://github.com/294coder/Efficient-MIF.
Multispectral pansharpening; multispectral and hyperspectral image fusion; Mamba; local-enhanced network
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3680905ccs: Computing methodologies Reconstruction
@@ 1. Introduction
In addressing the first question, two pioneering works stand out: Vim (Zhu et al., 2024), which introduced a bi-directional vision Mamba block tailored for image processing, and VMamba (Liu et al., 2024a), which proposed a scanning mechanism for images to tackle the causality issue inherent in Mamba’s unsuitability for vision tasks. As for the rest of the issues, there has been a lack of exploration of Mamba in the image fusion field. In this paper, we propose two bespoke techniques named local-enhanced vision Mamba (LEVM) block and state sharing to address issues 2) and 3), as its conception shown in Fig. 2, respectively enhancing Mamba’s perception of local information and fully exploiting spatial and spectral information in SSM’s state.
The proposed method is more suitable for image fusion tasks with two advantages to the proposed methods. The first advantage is that the proposed LEVM block can represent local and global spatial information. Local information brings rich spatial details, while global information injects long-range pixel dependencies, thus enhancing the detail recovery in the fusion process. The second advantage is that the state sharing method allows information to be shared between layers in the adjacent flow and skip-connected flow. Furthermore, we have also proposed the spatial-spectral learning (S2L) of SSM for the state sharing method. The S2L maps the input image to the state space for spatial representation, and considers the state space as the basis for spectral representation, thereby achieving interaction between spatial and spectral information. To the best of our knowledge, there have been no improvements made to vision Mamba from the perspective of spatial and spectral learning specifically for image fusion tasks. Finally, we propose the LE-Mamba, which is constructed based on the U-Net (Ronneberger et al., 2015a; Cao et al., 2024a, b) baseline with the inclusion of the LEVM block and the state sharing technique. LE-Mamba can achieve superior fusion performance.
@@ 2.2. Deep Learning Methods for Image Fusion
For the multispectral pansharpening task, many CNN-based methods have emerged and obtained promising fused images, including DiCNN (He et al., 2019), PanNet (Yang et al., 2017), and FusionNet (Deng et al., 2020), FeINFN (Liang et al., 2024). However, the local feature representation obtained by CNNs hinders better fusion results. For the multispectral and hyperspectral image fusion, there are also many outstanding DL-based works including InvFormer (Zhou et al., 2022b) and MiMO-SST (Fang et al., 2024). However, these methods bear a significant computational burden due to the quadratic complexity of the Transformer. Despite many efforts to mitigate this issue (Han et al., 2023; Wang et al., 2021; Liu et al., 2021), these methods usually lead to a performance drop.
@@ 4.1. Overall Architecture
In some high-level tasks (e.g., classification (Liu et al., 2024a; Zhu et al., 2024) and segmentation (Liu et al., 2024b; Ma et al., 2024; Ruan and Xiang, 2024; Liu et al., 2024a)), architectures similar to Meta-formers (Yu et al., 2022) are commonly used, where plain backbones can offer lower parameter counts, smaller computational loads, and lower latency. However, for the image fusion task, we opted for a multi-scale architecture akin to U-Net (Ronneberger et al., 2015b), which has been proven effective (Cao et al., 2024a, b). The designed network is illustrated in Fig. 2(d), featuring an encoder-decoder architecture. The input comprises LRMS and PAN images, while the output is fused images. The encoder and decoder are composed of multiple LEVM layers, with each LEVM layer comprising several LEVM blocks.
@@ 5.4.3. LEVM Outperforms Previous Attentions
To compare the proposed LEVM with previously common Attention mechanisms, we select common attention and its variants including traditional Attention (Dosovitskiy et al., 2021), PVT Attention (Wang et al., 2021), Swin Attention (Liu et al., 2021), and Linear Attention (Han et al., 2023). We replace the LEVM block with the corresponding Attention block (including FFN) and train them uniformly on the WV3 dataset until convergence. Their fusion performance is shown in Tab. 6.
It can be seen that the traditional Attention performs the worst, which is because using Attention in the first layer of the network occupies an extremely large GPU memory, leading to the inability to train. Therefore, most methods based on traditional Attention perform patch embedding in the first layer to downsample the image size, which is disastrous for tasks like image fusion that require low-level information. The performance of Linear Attention is also relatively poor, which is caused by the approximation errors introduced by its approximation of the Attention mechanism. Next are PVT Attention and Swin Attention, both of which impose restrictions on Attention. PVT downsamples both the keys and values, resulting in information loss, while Swin limits the global information to the local regions. The proposed LEVM outperforms previous attention operations.
Table 6. Results of different attention types on WV3 dataset.
@@
5.4.4. Enlarging SSM State Capacity Does not Solve Issue 2)
@@ 6. Conclusion
In conclusion, this paper introduces LE-Mamba, a multi-scale network architecture for image fusion. The proposed model addresses limitations in sequence modeling by implementing a local-enhanced vision Mamba block and state sharing technique. These innovations improve the network’s ability to capture both local and global information, enhancing spatial and spectral details in fused images. Experimental results demonstrate LE-Mamba’s state-of-the-art performance across various image fusion datasets.
7. Acknowledgement
This work is supported by the National Natural Science Foundation of China under Grants 12271083, 12171072.